# Banana Trees for the Persistence in Time Series Experimentally

### v1

@@ Abstract
In numerous fields, dynamic time series data require continuous updates, necessitating efficient data processing techniques for accurate analysis. This paper examines the banana tree data structure, specifically designed to efficiently maintain persistent homology—a multi-scale topological descriptor—for dynamically changing time series data. We implement this data structure and conduct an experimental study to assess its properties and runtime for update operations. Our findings indicate that banana trees are highly effective with unbiased random data, outperforming state-of-the-art static algorithms in these scenarios. Additionally, our results show that real-world time series share structural properties with unbiased random walks, suggesting potential practical utility for our implementation.
@@ 1 Introduction
  • Real-world data: through analysis of three specific datasets, we illustrate that real-world data can exhibit structural properties similar to those of unbiased random walks. Preliminary evidence suggests potential for the broad applicability of banana trees in practical scenarios (see Table 3).
In software for topological data analysis, tools like the Gudhi library [17], Dionysus [18], and Ripser [3] are established for static datasets but require complete recomputation for updates, making them less suited for dynamic contexts. We aim for our implementation to set a new standard in the topological processing of dynamic time series.
@@
Section 2 introduces persistent homology tailored to time series data and gives an overview of the banana trees data structure. Section 3 evaluates the performance of banana trees through experiments involving time series generated from random walks, with and without a bias. Section 4 explores three distinct types of input time series to assess the performance of banana trees: worst-case scenarios, quasi-periodic signals, and real-world data. Section 5 concludes the paper with a summary of our findings and their implications.
@@ 2.1 Persistent Homology of Time Series
The sublevel set of $f$ at $t\in\mathbb{R}$, denoted $f_{t}=f^{-1}(-\infty,t]$, are the points $x\in[0,n-1]$ that satisfy $f(x)\leq t$. When $t$ passes the value of a minimum from below, then the number of connected components of $f_{t}$ increases by $1$, and if $t$ passes the value of a maximum from below, the number of connected components decreases by $1$, unless the maximum is a endpoint, in which case the number does not change. Symmetrically, we call $f^{t}=f^{-1}[t,\infty)$ the superlevel set of $f$ at $t$. Observe that $f^{t}$ is the sublevel set of $-f$ at $-t$, and that the minima and maxima of $-f$ are the maxima and minima of $f$. Persistent homology tracks the evolution of the connected components while the sublevel set of $f$ grows, and formally defines when a component is born and when it dies. Complementing this with the same information for the superlevel sets of $f$, we get what is formally referred to as extended persistent homology; see [8] for details. It is best constructed in two phases:
@@ 2.2 Banana Trees
A time series represented by a piecewise linear function, $f$, is stored in two banana trees together with a linked list and two standard dictionaries. We call the first banana tree the up-tree of $f$ because it corresponds to Phase One of the persistence diagram construction. The second banana tree is the down-tree of $f$, which is the up-tree of $-f$. Because of this symmetry, it suffices to describe the up-tree of $f$. Its leaves are the minima of $f$, and its internal nodes are the maxima of $f$, with the exception of endpoint maxima, if they exist, since they are not critical in Phase One. Each point in the ordinary persistence diagram is a min-max pair, $(a,b)$, and it is represented by a banana that connects the leaf $a$ to the internal node $b$ with two parallel trails. The mid-trail is a doubly-linked list connecting $a$ to $b$ with the maxima that span windows nested in the mid-panel of the window of $a,b$, and the in-trail is symmetrically defined; see the right panel in Figure 1. The tree structure arises because the maxima belong to the similarly defined bananas of the nested windows. Each trail is sorted in the order of value but also in the order of location along the time series. For example, the mid-trail of the big banana in Figure 1 stores $a,s,q,b$, which montonically increases in value and monotonically decreases in location. Compare this with the in-trail storing $a,v,b$, which monotonically increases in value and in location after we discard $b$. Indeed, we think of $b$ as part of the min-trail while only connecting the in-trail to the rest of the tree without properly belonging to it. The global minimum of $f$ (not shown in Figure 1) is special because it is not paired in Phase One. Indeed, it is the extra leaf whose path upward does not end at an internal node. We therefore add a special root as the parent of the root, connected to the global minimum by the root banana. While there is only one root banana, there are possibly many leaf bananas, which are the bananas with two empty trails (beside the minimum and maximum they connect). Indeed, we call the number of nodes between the minimum and maximum the length of the trail, which for trails of leaf bananas is necessarily zero. Another important concept is the nesting depth of a banana, which is the number of windows that contain the corresponding window. The root banana has nesting depth zero, and all other bananas have positive nesting depth.
@@
Each update to a time series stored in a banana tree reduce to a sequence of elementary operations, which we name by their impact on the function, $f$. An interchange happens when two maxima or two minima swap the order of their values. Unless their locations are near each other, there is a good chance that an interchange does not have any effect on the banana tree. Otherwise, it is akin a rotation in a binary search tree, if the interchange is between two maxima, and akin a local rearrangement of the path decomposition, if the interchange is between two minima. A cancellation removes a min-max pair or, equivalently, the corresponding leaf banana, an anti-cancellation is the inverse of a cancellation, and a slide swaps the status of a critical item with that of a neighboring non-critical item. The latter three operations appear only a constant number of times whenever we adjust a value. Nevertheless, anti-cancellations present challenges to fast implementation as we have to find out where it is to happen, and this search is not supported by any special purpose data structure. On the other hand, interchanges may happen in sequence, so it is important to charge each to a change in the persistence diagram. We refer to [9] and Appendices A and B for details.
@@ 3.1 The Experimental Set-up
We write $\gamma\sim\mathcal{N}(\mu,\sigma)$ for a normally distributed random variable with mean $\mu$ and standard deviation $\sigma$. Given $\mu,\sigma$, a random walk of length $n$ is a real-valued function $r=r_{\mu,\sigma}$ defined by
[…]
@@
To evaluate the performance of local updates, we change the value of a single item, and since it depends on the amount, we fix parameters $\Delta\in\mathbb{R}$ and $k\in\mathbb{N}$ and change the value of the $i$-th item from $r(i)$ to $r(i)+\delta_{j}$, with $\delta_{j}=\frac{j}{k}\Delta$ for $-k\leq j\leq k$. For the experiments, we pick $\Delta=5.0$ or $50.0$ and $k=10$, while doing the update on a random walk of length between $10^{2}$ and $10^{6}$ generated for $\mu\in[-1.1]$ and $\sigma=1$. Each experiment is repeated one hundred times and averages as well as deviations from the average are reported. For each experiment, we also measure the time to run Gudhi on the walk after the value change. This provides the appropriate reference time, since Gudhi needs to recompute the persistence diagram from scratch after each update.
To evaluate the performance of topological updates (split or glue), we pick a fraction, $c\in(0,1)$, generate a random walk of length $n$, and then cut it into a left walk of length $\lfloor{cn}\rfloor$ and a right walk of length $\lceil{(1-c)n}\rceil$. To split, we first construct the banana tree of the entire random walk, which we then cut into two banana trees, one for the left and the other for the right walk. To glue, we first construct the banana trees of the left and right walks, which we then combine to a single banana tree for the walk of length $n$. Doing this for fractions $c\in\{0.1,0.3,0.5,0.7,0.9\}$, we compare the time to split with the time used by Gudhi to construct the persistence diagrams of the left and right walks, and the time to glue with the time used by Gudhi to construct the persistence diagram of the entire random walk.
@@ 3.2 Structural Properties
[…]
Figure 2: Left: the average fraction of items that are critical as a function of the bias. Middle: the average nesting depth of leaf bananas in banana trees of unbiased random walks with $n$ items. The ribbon extends from the minimum to the maximum observed value for each $n$; the dots mark the mean. The red line is the graph of a constant times $\log n$ obtained by linear regression. Right: the average length of in-trails (orange dots) and mid-trails (blue triangles) in banana trees of random walks with bias $\mu$, averaged over all input sizes.
The middle panel of Figure 2 shows the average nesting depth of a leaf banana for an unbiased random walk, which appears to scale like the logarithm of the number of items. In the unbiased case, even the maximum nesting depth seems to be small (about $13$ on average for a random walk of length $2\cdot 10^{6}$), and scale logarithmically in the number of items. With increasing bias, the maximum nesting depth decreases, to about $4$ at $\mu=\pm 1$. This is because in the biased case, there are fewer critical items but also the bananas tend to arrange in parallel rather than inside each other. We observe that the nesting depth is about the same on average in up-trees and down-trees, both for biased and unbiased random walks, so our results hold for both trees.
Regarding trails, we observe that the length of the in-trails in the up-tree and the mid-trails in the down-tree are equal on average. By symmetry, the same holds for the mid-trails in the up-tree and the in-trails in the down-tree. Let us therefore focus on the up-tree. The right panel in Figure 2 shows how the average length of in- and mid-trails depends on the bias. In the unbiased case, both types of trails have the same length on average. For positive $\mu$, the mid-trails tend to be longer than the in-trails, while the reverse happens for negative $\mu$. The main cause for this trend is the longest trail, which starts to dominate the others in length as the bias increases. Assuming $\mu>0$, the global minimum and maximum tend to lie to the left and right of the middle, respectively, with the majority of the items between them. By convention, the connecting in-trail and mid-trail contain the items to the left and right of the global minimum, respectively. This explains why the mid-trail dominates in this case, and why the in-trail dominates when $\mu<0$. Indeed, for $\mu=\pm 1$, we observe about $95\%$ of the internal nodes (the local maxima) on the longest trail.
@@ 3.4 Topological Maintenance
The advantage of the banana tree over Gudhi diminishes when the random walks get progressively more biased. The reason is again that the larger the bias, the larger the fraction of internals nodes in the longest trail of the banana tree. Splitting the corresponding banana requires the resetting of many pointers stored in nodes along this trail, which in some cases costs time proportional to the number of critical items, which we indeed observe in our experiments. Unfortunately, this more than compensates for the low nesting depth, which guarantees that only a small number of bananas need to be split. The upper half of Table 1 shows the running time and the speedup over Gudhi for cutting a random walk in half.
Table 1: Upper half: the average time for cutting a random walk of length $n=10^{6}$ in half, and the speedup over Gudhi. Lower half; the average time for concatenating two random walks of length $n=10^{6}/2$ each, and the speedup over Gudhi. All times are measured in micro-seconds.
@@
We finally address the reverse operation, that of concatenating two random walks or, correspondingly, of gluing two banana trees. The curves in the right panel of Figure 3 are similar to those in the middle panel, which suggests that the performance is similar to that of cutting, which is confirmed by the numbers in Table 1. For $n\leq 10^{4}$, gluing appears to be slightly fast than splitting, but the difference is less clear already for $n=10^{6}$.
@@ 4 Special Time Series
This section considers three types of time series: worst-case examples to expose when banana trees underperform, quasi-periodic data to illustrate how banana trees reflect periodicity, and real-world data to demonstrate that banana trees are not just theory. While banana trees struggle for data of the first type, they mirror the performance observed in unbiased random walks for the latter two types.
@@ 4.1 Worst-case Examples
Consider first the operation that increases the value of the marked item in the left panel of Figure 4. When its value passes that of its left neighbor, an anti-cancellation adds the banana they span to the banana tree. As we increase the value further, the marked item interchanges with the maxima to its right, each time moving up one level within the sequence of windows within which its window is nested. When the marked item finally becomes the global maximum, it will have interchanged with all other maxima, which takes $\Theta(n)$ time. For $n=10^{2}$ and $n=10^{6}$, we measured $3.4\mu{\rm s}$ and $4.7{\rm m}{\rm s}$ on average, respectively. This corresponds to a median speedup of $0.05$ and $0.02$ if compared with Gudhi, which is really a slowdown by a factor of $20$ and $50$, respectively.
@@
$\lambda=0$ $10^{-4}$ $10^{-2}$ $10^{0}$ $\lambda=0$ $10^{-4}$ $10^{-2}$ $10^{0}$
$n=10^{2}$ 0.04 0.04 0.04 0.30 0.07 0.07 0.09 0.55
$n=10^{4}$ 0.02 0.02 4.60 10.00 0.05 0.06 7.80 18.00
$n=10^{6}$ 0.04 3.60 41.00 100.00 0.05 3.70 55.00 102.00
Consider second the operation that cuts the time series in the right panel of Figure 4 in the middle, which is marked by the vertical dashed line. We observe an average running time between $14\mu{\rm s}$ at $n=10^{2}$ and $98{\rm m}{\rm s}$ at $n=10^{6}$. Compare this with an average running time of $10\mu{\rm s}$ at $n=10^{2}$ and $80{\rm m}{\rm s}$ at $n=10^{6}$ for concatenating the two lists back to the original length. This comparison agrees with the earlier observation that gluing banana trees is slightly faster than splitting them. Table 2 lists the speedup if compared to Gudhi, which for a number less than $1$ is really a slowdown. To create a more informative experiment, we interpolate between these time series and unbiased walks, with drastic improvements even for small $\lambda>0$; see Table 2. Indeed, the noise quickly flattens the banana tree by decreasing the length of trails, which explains the improvement.

### v2

@@ Abstract
In numerous fields, dynamic time series data require continuous updates, necessitating efficient data processing techniques for accurate analysis. This paper examines the banana tree data structure, specifically designed to efficiently maintain the multi-scale topological descriptor commonly known as persistent homology for dynamically changing time series data. We implement this data structure and conduct an experimental study to assess its properties and runtime for update operations. Our findings indicate that banana trees are highly effective with unbiased random data, outperforming state-of-the-art static algorithms in these scenarios. Additionally, our results show that real-world time series share structural properties with unbiased random walks, suggesting potential practical utility for our implementation.
keywords:
persistent homology, time series, data structures, computational experiments.
@@ 1 Introduction
  • Real-world data: through analysis of three specific datasets, we illustrate that real-world data sometimes exhibits structural properties similar to those of unbiased random walks. Preliminary evidence suggests potential for the broad applicability of banana trees in practical scenarios (see Table 3).
In software for topological data analysis, tools like the Gudhi library [16], Dionysus [17, 18], and Ripser [2] are established for static datasets but require complete recomputation for updates, making them less suited in dynamic settings. We aim for our implementation to set a new standard in the topological processing of dynamic time series.
@@
Section 2 introduces persistent homology tailored to time series data and gives an overview of the banana trees data structure. Section 3 evaluates the performance of banana trees through experiments involving time series generated from random walks, with and without a bias. Section 4 explores three distinct types of input time series to assess the performance of banana trees: worst-case scenarios, quasi-periodic signals, and real-world data. Section 5 concludes the paper with a summary of our findings.
@@ 2.1 Persistent Homology of Time Series
The sublevel set of $f$ at $t\in\mathbb{R}$, denoted $f_{t}=f^{-1}(-\infty,t]$, are the points $x\in[0,n-1]$ that satisfy $f(x)\leq t$. When $t$ passes the value of a minimum from below, then the number of connected components of $f_{t}$ increases by $1$, and if $t$ passes the value of a maximum from below, the number of connected components decreases by $1$, unless the maximum is an endpoint, in which case the number does not change. Symmetrically, we call $f^{t}=f^{-1}[t,\infty)$ the superlevel set of $f$ at $t$. Observe that $f^{t}$ is the sublevel set of $-f$ at $-t$, and that the minima and maxima of $-f$ are the maxima and minima of $f$. Persistent homology tracks the evolution of the connected components while the sublevel set of $f$ grows, and formally defines when a component is born and when it dies. Complementing this with the same information for the superlevel sets of $f$, we get what is formally referred to as extended persistent homology; see [6] for details. It is best constructed in two phases:
@@ 2.2 Banana Trees
A time series represented by a piecewise linear function, $f$, is stored in two banana trees together with a linked list and two standard dictionaries. We call the first banana tree the up-tree of $f$ because it corresponds to Phase One of the persistence diagram construction. The second banana tree is the down-tree of $f$, which is the up-tree of $-f$. Because of this symmetry, it suffices to describe the up-tree of $f$. Its leaves are the minima of $f$, and its internal nodes are the maxima of $f$, with the exception of endpoint maxima, if they exist, since they are not critical in Phase One. Each point in the ordinary persistence diagram is a min-max pair, $(a,b)$, and it is represented by a banana that connects the leaf $a$ to the internal node $b$ with two parallel trails. The mid-trail is a doubly-linked list connecting $a$ to $b$ with the maxima that span windows nested in the mid-panel of the window of $a,b$, and the in-trail is symmetrically defined; see the right panel in Figure 1. The tree structure arises because the maxima belong to the similarly defined bananas of the nested windows. Each trail is sorted in the order of value but also in the order of location along the time series. For example, the mid-trail of the big banana in Figure 1 stores $a,s,q,b$, which monotonically increases in value and monotonically decreases in location. Compare this with the in-trail storing $a,v,b$, which monotonically increases in value and in location after we discard $b$. Indeed, we think of $b$ as part of the mid-trail while only connecting the in-trail to the rest of the tree without properly belonging to it. The global minimum of $f$ (not shown in Figure 1) is special because it is not paired in Phase One. Indeed, it is the extra leaf whose path upward does not end at an internal node. We therefore add a special root as the parent of the root, connected to the global minimum by the root banana. While there is only one root banana, there are possibly many leaf bananas, which are the bananas with two empty trails (beside the minimum and maximum they connect). Indeed, we call the number of nodes strictly between the minimum and maximum the length of the trail, which for trails of leaf bananas is necessarily zero. Another important concept is the nesting depth of a banana, which is the number of windows that contain the corresponding window. The root banana has nesting depth zero, and all other bananas have positive nesting depth.
@@
Each update to a time series stored in a banana tree reduces to a sequence of elementary operations, which we name by their impact on the function, $f$. An interchange happens when two maxima or two minima swap the order of their values. Unless their locations are near each other, there is a good chance that an interchange does not have any effect on the banana tree. Otherwise, it is akin a rotation in a binary search tree, if the interchange is between two maxima, and akin a local rearrangement of the path decomposition, if the interchange is between two minima. A cancellation removes a min-max pair or, equivalently, the corresponding leaf banana, an anti-cancellation is the inverse of a cancellation, and a slide swaps the type of a critical item with that of a neighboring non-critical item. The latter three operations appear only a constant number of times whenever we adjust a value. Nevertheless, anti-cancellations present challenges to fast implementation as we have to find out where the new banana is to be inserted, and this search is not supported by any special purpose data structure. On the other hand, multiple interchanges may happen in sequence, so it is important to charge the time for processing each to a change in the persistence diagram. We refer to [8] and Appendix A for details.
@@ 3 Experimental Results for Random Walks
We implemented the banana trees data structure in C++20, using the AVL trees and splay trees provided by boost.intrusive [14]. The former offer good performance in general, the latter allow for simple splitting and joining. Hence, we use splay trees when performing topological operations and AVL trees everywhere else. The experiments were run on a machine with two AMD EPYC 9534 64 core CPUs and 1.5 terabyte of main memory.
@@ 3.1 The Experimental Set-up
We write $\gamma\sim\mathcal{N}(\mu,\sigma)$ for a normally distributed random variable with mean $\mu$ and standard deviation $\sigma$. Given $\mu,\sigma$, a random walk of length $n$ is a function, $r=r_{\mu,\sigma}$, from the first $n$ non-negative integers to the reals, defined by
$r(0)=0\mbox{\rm\leavevmode\nobreak\ \leavevmode\nobreak\ and\leavevmode\nobreak\ \leavevmode\nobreak\ }r(i)=r(i-1)+\gamma_{i}\mbox{\rm\leavevmode\nobreak\ \leavevmode\nobreak\ for\leavevmode\nobreak\ }1\leq i<n,$ (1)
@@
To evaluate the performance of local updates, we change the value of a single item, and since it depends on the amount, we fix parameters $\Delta\in\mathbb{R}$ and $k\in\mathbb{N}$ and do $12k+$ 1 parallel updates by changing the value of the $i$-th item from $r(i)$ to $r(i)+\delta_{j}$, with $\delta_{j}=\frac{j}{k}\Delta$ for all integers $j$ with $-k\leq j\leq k$. For the experiments, we pick $\Delta=5.0$ or $50.0$ and $k=10$, while doing the update on a random walk of length between $10^{2}$ and $10^{6}$ generated for $\mu\in[-1,1]$ and $\sigma=1$. Each experiment is repeated one hundred times and averages as well as deviations from the average are reported. For each experiment, we also measure the time to run Gudhi on the walk after the value change. This provides the appropriate reference time, since Gudhi needs to recompute the persistence diagram from scratch after each update.
To evaluate the performance of topological updates, we pick a fraction, $c\in(0,1)$, generate a random walk of length $n$, and then cut it into a left walk containing the first $\lfloor{cn}\rfloor$ items and a right walk containing the remaining $\lceil{(1-c)n}\rceil$ items. There are two such updates, which we now define. To split, we first construct the banana tree of the entire random walk, which we then cut into two banana trees, one for the left and the other for the right walk. To glue, we first construct the banana trees of the left and right walks, which we then combine to a single banana tree for the walk of length $n$. Doing this for fractions $c\in\{0.1,0.3,0.5,0.7,0.9\}$, we compare the time to split with the time used by Gudhi to construct the persistence diagrams of the left and right walks, and the time to glue with the time used by Gudhi to construct the persistence diagram of the entire random walk.
@@ 3.2 Structural Properties
Figure 2: Left: the average fraction of items that are critical as a function of the bias $\mu$. Middle: the average nesting depth of leaf bananas in banana trees of unbiased random walks with $n$ items. The ribbon extends from the minimum to the maximum observed value for each $n$; the dots mark the mean. The red line is the graph of a constant times $\log n$ obtained by linear regression. Right: the maximum nesting depth at $n=10^{6}$ items as a function of the bias $\mu$.
Figure 3: Left: the average length of in-trails (orange dots) and mid-trails (blue triangles) in banana trees of random walks with bias $\mu$, averaged over all input sizes. Right: the fraction of nodes on the longest in-trail (orange dots) and longest mid-trails (blue triangles) in banana trees of random walks with bias $\mu$, averaged over all input sized.
The middle panel of Figure 2 shows the average nesting depth of a leaf banana for an unbiased random walk, which appears to scale like the logarithm of the number of items. In the unbiased case, even the maximum nesting depth seems to be small (about $13$ on average for a random walk of length $2\cdot 10^{6}$), and scale logarithmically in the number of items. With increasing bias, the maximum nesting depth decreases, to about $4$ at $\mu=\pm 1$; see the right panel of Figure 2. This is because in the biased case, there are fewer critical items but also the bananas tend to arrange in parallel rather than inside each other. We observe that the nesting depth is about the same on average in up-trees and down-trees, both for biased and unbiased random walks, so our results hold for both trees.
For topological reasons, the mid-trails of the up-tree are the mid-trails of the down-tree, except that they are upside down. In contrast, the in-trails of the two trees are not directly related because the in-panels in one direction become the out-panels in the other. Notwithstanding, we observe that the length of the in-trails in the up-tree and the mid-trails in the down tree are very similar on the average, and by symmetry the same holds for the mid-trails in the up-tree and the in-trails in the down-tree. These two facts do not contradict each other, because for biased data most of the critical items belong to the root banana, which does not adhere to the topological constraints just mentioned. Let us therefore focus on the up-tree. The left panel in Figure 3 shows how the average length of in- and mid-trails depends on the bias. In the unbiased case, both types of trails have the same length on average. For positive $\mu$, the mid-trails tend to be longer than the in-trails, while the reverse happens for negative $\mu$. The main cause for this trend is the longest trail, which starts to dominate the others in length as the bias increases; see the right panel in Figure 3. Assuming $\mu>0$, the global minimum and maximum tend to lie to the left and right of the middle, respectively, with the majority of the items between them. By convention, the connecting in-trail and mid-trail contain the items to the left and right of the global minimum, respectively. This explains why the mid-trail dominates in this case, and why the in-trail dominates when $\mu<0$. Indeed, for $\mu=\pm 1$, we observe about $95\%$ of the internal nodes (the local maxima) on the longest trail.
@@ 3.4 Topological Maintenance
The advantage of the banana tree over Gudhi diminishes when the random walks get progressively more biased. The reason is again that the larger the bias, the larger the fraction of internal nodes in the longest trail of the banana tree. Splitting the corresponding banana requires the resetting of many pointers stored in nodes along this trail, which in some cases costs time proportional to the number of critical items, which we indeed observe in our experiments. Unfortunately, this more than compensates for the low nesting depth, which guarantees that only a small number of bananas need to be split. The upper half of Table 1 shows the running time and the speedup over Gudhi for cutting a random walk in half.
Table 1: Upper half: the average time for cutting a random walk of length $n=10^{6}$ in half, and the speedup over Gudhi. Lower half; the average time for concatenating two random walks of length $n=10^{6}/2$ each, and the speedup over Gudhi.
@@
We finally address the reverse operation, that of concatenating two random walks or, correspondingly, of gluing two banana trees. The curves in the right panel of Figure 4 are similar to those in the middle panel, which suggests that the performance is similar to that of cutting, which is confirmed by the numbers in Table 1. For $n\leq 10^{4}$, gluing appears to be slightly faster than splitting, but the difference is less clear already for $n=10^{6}$.
@@ 4 Special Time Series
This section considers three types of time series: worst-case examples to expose when banana trees under-perform, quasi-periodic data to illustrate how banana trees reflect periodicity, and real-world data to demonstrate that banana trees are not just theory. While banana trees struggle for data of the first type, they mirror the performance observed in unbiased random walks for the latter two types.
@@ 4.1 Worst-Case Examples
Consider first the operation that increases the value of the marked item in the left panel of Figure 5. Initially, we have a linear sequence of double-panel windows, each nested inside the next. When the value of the marked item passes that of its left neighbor, an anti-cancellation adds the banana they span to the banana tree, and its window is nested inside the others. As we increase the value further, the marked item interchanges with the maxima to its right, each time moving up one level within the sequence of windows. When the marked item finally becomes the global maximum, it will have interchanged with all other maxima, which takes $\Theta(n)$ time. For $n=10^{2}$ and $n=10^{6}$, we measured $3.4\mu{\rm s}$ and $4.7{\rm m}{\rm s}$ on average, respectively. This corresponds to a median speedup by a factor of $0.05$ and $0.02$ if compared with Gudhi, which is really a slowdown by a factor of $20$ and $50$, respectively.
@@
$\lambda=0.0$ $0.0001$ $0.01$ $1.0$ $\lambda=0.0$ $0.0001$ $0.01$ $1.0$
$n=10^{2}$ 0.06 0.06 0.04 0.35 0.40 0.44 0.68 2.38
$n=10^{4}$ 0.04 0.03 5.81 11.20 0.46 0.83 33.6 41.40
$n=10^{6}$ 0.09 4.71 54.70 111.00 0.46 81.0 255 397.00
Consider second the operation that cuts the time series in the right panel of Figure 5 in the middle, which is marked by the vertical dashed line. We observe an average running time between $10\mu{\rm s}$ at $n=10^{2}$ and $72{\rm m}{\rm s}$ at $n=10^{6}$. Compare this with an average running time of $7\mu{\rm s}$ at $n=10^{2}$ and $48{\rm m}{\rm s}$ at $n=10^{6}$ for concatenating the two lists back to the original length. This comparison agrees with the earlier observation that gluing banana trees is slightly faster than splitting them. Table 2 lists the speedup if compared to Gudhi, which for a factor less than $1$ is really a slowdown. To create a more informative experiment, we interpolate between these time series and unbiased walks, with drastic improvements even for small $\lambda>0$; see Table 2. Indeed, the noise quickly flattens the banana tree by decreasing the length of trails, which explains the improvement.